OpenStudy (abbles):

Identify the extrema for each function. Classify each as a relative (local) or absolute (global) maximum or minimum value.

OpenStudy (abbles):

\[f(x) = 4x^3 - 12x^2 + 15\]

OpenStudy (sshayer):

\[f \prime \left( x \right)=12x^2-24x\] \[f \prime \left( x \right)=0~gives~12x^2-24x=0,12x(x-2)=0,x=0,2\] \[f \prime \prime \left( x \right)=24x-12\] \[f \prime \prime \left( 0 \right)=0-12=-12<0\] There is a local maxima at x=0 f(0)=? when x=2, f"(2)=24*2-12=36>0 hence there is a relative minima and minimum value is f(2)=?

OpenStudy (sshayer):

correction

OpenStudy (abbles):

So I set the function equal to 0 and solve to find the local maximum?

OpenStudy (sshayer):

\[f \prime \prime (x)=24x-24\] then proceed further as i have done.

OpenStudy (sshayer):

if you want local maxima then put x=0 if local minima then put x=2

OpenStudy (sshayer):

f(0)=4*0^3-12*0^2-15=? f(2)=?

OpenStudy (abbles):

Okay. f(x)=4x3−12x2+15 f(0) = -15 -15 would be the local maximum? how do I know it's the "local" maximum and not the absolute maximum? f(x)=4x3−12x2+15 f(2) = 32 - 48 + 15 f(2) = -1 The local (or absolute?) minimum equals -1. Right?

OpenStudy (abbles):

Is the formula always f(2) = the local minimum, or just for this example?

OpenStudy (sshayer):

sorry it is +15 not -15

OpenStudy (abbles):

Right.. :P

OpenStudy (sshayer):

no here we found critical values by putting f"(x)=0 they are the values

OpenStudy (abbles):

?

OpenStudy (sshayer):

we have to check with these critical values.

OpenStudy (peachpi):

@Abbles are you a calculus student?

OpenStudy (abbles):

precalculus

OpenStudy (peachpi):

but do you understand what @sshayer is doing with the derivatives?

OpenStudy (abbles):

Um... what are derivatives? :P

OpenStudy (abbles):

Also, what are "critical values"?

OpenStudy (abbles):

To be clear... @sshayer was saying that to find the local maximum of a function, evaluate f(0) and to find the local minimum, evaluate f(2). Is that true for all functions? what about the absolute maximum/min?

jimthompson5910 (jim_thompson5910):

@Abbles it sounds like you haven't learned derivatives yet since you're in pre-calculus. So I'd use a TI calculator (eg: TI83) or desmos or geogebra to graph the equation y = 4x^3 - 12x^2 + 15. Then use the min max feature to find the extrema. https://www.desmos.com/calculator/1eheltseuy The nice thing about desmos is that you simply click the extrema points to have the coordinates show up. See attached.

OpenStudy (abbles):

Alrighty, got my TI-84 ready. Is there no way to find the min/max without a calculator?

jimthompson5910 (jim_thompson5910):

`Alrighty, got my TI-84 ready. Is there no way to find the min/max without a calculator?` the only way without a calculator is to use calculus but that will come in a later class anyways, graph the function using your TI-84 by typing in the equation in the y1 slot (after hitting the \(\Large y=\) key in the top left corner). Then hit the \(\large \text{GRAPH}\) key at the top right corner to see the graph. Adjust the window so you see something similar to what I posted from the desmos graph. Once the graph is all set up, hit the blue \(\large \text{2ND}\) key and then hit the \(\large \text{TRACE}\) key (next to the GRAPH key). Scroll down to 3:minimum and that will help find the local min. You'll have to provide the left boundary, the right boundary, and a guess somewhere in the middle. Repeat these steps, but do 4:maximum instead to get the local max. Again you'll need the left boundary, the right boundary, and a guess somewhere in the middle.

OpenStudy (abbles):

After I plugged in the function and clicked GRAPH, it's pulling up ERROR: INVALID DIMENSION

OpenStudy (abbles):

:/ any idea what's wrong? the calculator is fairly new and I haven't fooled around with it much.

jimthompson5910 (jim_thompson5910):

you typed this in right? see attached

OpenStudy (abbles):

Yes pretty much... the model is actually a TI-84 plus ce... but same concept

jimthompson5910 (jim_thompson5910):

what do you have for your window?

OpenStudy (abbles):

Xmin = -10 and Xmax = 10 ? Not sure what all these mean... Xscl = 1 Yscl = 1 Xres=1 etc.

jimthompson5910 (jim_thompson5910):

this is what I have.

jimthompson5910 (jim_thompson5910):

go ahead and type in what I have so we have the same window

OpenStudy (abbles):

Done. Still an error though. It says: Check 1<dim(list)<999. To set Plots0ff: 2nd STAT PLOT; Plots0ff Check 1<dim(matrix)<99. Check inverse of square matrix only.

OpenStudy (abbles):

I'm already taking one foreign language class right now lol.

jimthompson5910 (jim_thompson5910):

so you have your stats plot on right now?

OpenStudy (abbles):

Sorry, what? :P I'm a complete newb when it comes to graphing calculators

jimthompson5910 (jim_thompson5910):

when you hit the \(\Large y=\) key (top left corner), do you see Plot1, Plot2, or Plot3 highlighted at all?

jimthompson5910 (jim_thompson5910):

if possible please show me a snapshot of the calculator screen you are looking at

OpenStudy (abbles):

Yes, Plot1

jimthompson5910 (jim_thompson5910):

ok go up to that and turn that off

jimthompson5910 (jim_thompson5910):

hitting enter over it should turn it off

OpenStudy (abbles):

Perfect, it's working now. So how do I find the mins/maxs of the graph? It's not showing any numbers on the graph for some reason...

jimthompson5910 (jim_thompson5910):

hit the blue \(\large \text{2ND}\) key and then hit the \(\large \text{TRACE}\) key (next to the GRAPH key). Scroll down to 3:minimum and that will help find the local min. You'll have to provide the left boundary, the right boundary, and a guess somewhere in the middle.

jimthompson5910 (jim_thompson5910):

make sure that the left boundary is to the left of where you think the min is. You'll have to visually eyeball things a bit. Make sure the right boundary is to the right of the min

OpenStudy (abbles):

For the minimum I got X=2.0000015 and y=-1

jimthompson5910 (jim_thompson5910):

due to some rounding error, you got 2.0000015 for x I got 1.9999997 for x (see attached) it should be x = 2 but the calculator only calculates to so many decimal digits and there's a bit of error.

OpenStudy (abbles):

Maximum: X=1.4772781 and Y-1.7075478

jimthompson5910 (jim_thompson5910):

So the min is (2,-1)

OpenStudy (abbles):

Okay

jimthompson5910 (jim_thompson5910):

`Maximum: X=1.4772781 and Y-1.7075478` the local max is nowhere close to this

OpenStudy (abbles):

Let me try again

jimthompson5910 (jim_thompson5910):

I should clarify that the local min is (2,-1) it is not an absolute minimum

OpenStudy (abbles):

How do you determine if it's an absolute or local min/max?

OpenStudy (abbles):

Yep, for the max I got a perfect x = 0 and y = 15

jimthompson5910 (jim_thompson5910):

If you have a certain interval, then the absolute max is the highest point on the interval. If you don't define an interval, then the absolute max is the entire domain. No intervals are stated in this problem, so we look at the entire cubic function. It goes on forever upward as x gets bigger, so there is NO absolute max at all

jimthompson5910 (jim_thompson5910):

The same can be said in the other direction as well

jimthompson5910 (jim_thompson5910):

so we have local extrema and no global (absolute) extrema

OpenStudy (abbles):

Got it

OpenStudy (abbles):

The absolute extrema can only be found if the domain is specified?

OpenStudy (abbles):

Mind going through one more with me? 3x^3 + 3x^2 - 30x + 24

jimthompson5910 (jim_thompson5910):

Let's say you had this parabola http://tutorial.math.lamar.edu/Classes/Alg/Parabolas_files/image005.gif over the entire domain, the lowest point is at (1,2). This is the global/absolute min in this problem. Unless a different interval is stated, we use the domain as our interval

jimthompson5910 (jim_thompson5910):

With that parabola, there is global max if no interval is stated because the graph keeps going up forever.

jimthompson5910 (jim_thompson5910):

`Mind going through one more with me? 3x^3 + 3x^2 - 30x + 24` sure, tell me how far you get with this next one

OpenStudy (abbles):

Okay, for this one I just need the local max. x=-2 and y=72 sound about right?

jimthompson5910 (jim_thompson5910):

Is it asking to round to the nearest whole number?

OpenStudy (abbles):

No...

OpenStudy (abbles):

To the nearest thousandth

jimthompson5910 (jim_thompson5910):

This is what I get https://www.desmos.com/calculator/pwvx206d96 see attached

OpenStudy (abbles):

72.5 seem closer? how am I supposed to find an answer as accurate as the thousandth place?

jimthompson5910 (jim_thompson5910):

72.5 is what your TI84 is saying?

OpenStudy (abbles):

No, from looking at your graph... that's what it looks to be

jimthompson5910 (jim_thompson5910):

oh, I recommend you use your TI84 again or use desmos like I did. Look at the last attachment I posted

OpenStudy (abbles):

I got it wrong :/ I put 72.5; the correct answer was 72.577-72.578 Oh well. At least now I know how to do it. Thanks for your help :)

jimthompson5910 (jim_thompson5910):

it should be 72.578 hopefully you see the coordinates on that attachment?

OpenStudy (abbles):

Thanks. I don't get why my calculator was being so inaccurate :/ oh well..